it took 40 lb. to reseed 5,000 yd squared....... how many yd squared could be reseeded with 225 lb.?
Answer by stanbon(75887)

it took 40 lb. to reseed 5,000 yd squared....... how many yd squared could be reseeded with 225 lb.?
---------------
Use a proportion::
area/wt = area/wt
-------
area/225 = 5000/40
-------
area = 225(5000/40) = 225(125) = 28,125 sq yd
------------------
Cheers,
Stan H.
-----------------